16/3y = 12/5
The slash is a fraction symbol.
: 16/3y = 12/5
The slash is a fraction symbol.

Answer by checkley77(3639)

16/3y=12/5
IF THIS IS 16/(3Y)=12/5 THEN:
12*3Y=16*5
36Y=80
Y=80/36
Y=20/9 ANSWER.
HOWEVER, IF THIS IS 16Y/3=12/5 THEN:
16Y*5=3*12
80Y=36
Y=36/80
Y=9/20 ANSWER.